# D-tensor paraproducts and its caricatures

### v1

@@ 1 Introduction
$\mathbf{v}^{d-1}\coloneq\frac{\partial^{2}}{\partial t_{d}\partial t_{d-1}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-2}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))+\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{d-1}\partial t_{d-2}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))$
$+\ldots+\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{3}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{2}\partial t_{1}}(\delta^{t}(f))\in[0,1]^{d}\to\mathbb{R}$ (33)
$\mathbf{v}^{d}\coloneq\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))\in[0,1]^{d}\to\mathbb{R}$ (34)
@@
$A(f)=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}\int_{0}^{1}\int_{0}^{1}\ldots\int_{0}^{1}A(\delta^{t}(f))\mathbf{v}^{1}+A(\delta^{t}(f))\mathbf{v}^{2}+\ldots+$
$A^{m}(\delta^{t}(f))\mathbf{v}^{m}+\ldots+A^{d-1}(\delta^{t}(f))\mathbf{v}^{d-1}+A^{d}(\delta^{t}(f))\mathbf{v}^{d}dt_{1}dt_{2}\ldots dt_{d}$ (35)
We can now define the approximation, $\tilde{A}(f)$, by replacing the sequence $(\mathbf{v}^{1},\mathbf{v}^{2}\ldots,\mathbf{v}^{d})$ with $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ - each element in the latter sequence is comprised with a wavelet coefficient in the direction of the partial derivative for an element in the former sequence.
$\mathbf{\tilde{v}}^{1}\coloneq Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}}(f)\in[0,1]^{d}\to\mathbb{R}$ (36)
$\tilde{\mathbf{v}}^{2}\coloneq Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{2}}(f)Q^{j_{1}}(f)+Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{3}}(f)Q^{j_{2}}Q^{j_{1}}(f)+\ldots$
$+Q^{j_{d}}(f)Q^{j_{d-1}}Q^{j_{d-2}}\ldots Q^{j_{1}}(f)\in[0,1]^{d}\to\mathbb{R}$ (37)
@@
$\tilde{\mathbf{v}}^{m}\coloneq Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{m}}(f)Q^{j_{m-1}}(f)\ldots Q^{j_{1}}(f)$
@@
$\tilde{\mathbf{v}}^{d-1}\coloneq Q^{j_{d}}Q^{j_{d-1}}(f)Q^{j_{d-2}}(f)\ldots Q^{j_{1}}(f)+Q^{j_{d}}(f)Q^{j_{d-1}}Q^{j_{d-2}}(f)\ldots Q^{j_{1}}(f)$
$+\ldots+Q^{j_{d}}(f)Q^{j_{d-1}}(f))\ldots Q^{j_{3}}(f)Q^{j_{2}}Q^{j_{1}}(f)\in[0,1]^{d}\to\mathbb{R}$ (39)
$\tilde{\mathbf{v}}^{d}\coloneq Q^{j_{d}}(f)Q^{j_{d-1}}(f)\ldots Q^{j_{1}}(f)\in[0,1]^{d}\to\mathbb{R}$ (40)
Again, note the first, $\tilde{\mathbf{v}}^{1}$, and last, $\tilde{\mathbf{v}}^{d}$, operators have one term while the operators, $(\mathbf{v}^{2},\ldots,\mathbf{v}^{d-1})$, are comprised of a series each with d1d-1 terms. In this case, each term in the series contains the number of wavelet coefficients corresponding to the power of the operator (i.e. the term $\tilde{\mathbf{v}}^{m}$ has $m$ wavelet coefficients for each term in its series). Now the following linear combination of $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ is independent of $t$ and is used to construct the approximation, $\tilde{A}(f)$:
$\tilde{A}(f)\coloneq\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}A(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}+A^{\prime\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{2}+\ldots+A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m}$
@@
$=\sum_{j_{d}=0_{d}}^{N_{d}}\sum_{j_{d-1}=0_{d-1}}^{N_{d-1}}\ldots\sum_{j_{1}=0_{1}}^{N_{1}}(\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha})$ (41)
Computing the residual, $\Delta(A,f)\coloneq A(f)-\tilde{A}(f)$, yields:
@@
$=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}\int_{0}^{1}\int_{0}^{1}\ldots\int_{0}^{1}\sum_{m=1}^{d}(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})dt_{1}dt_{2}\ldots dt_{d}$ (42)
We characterize the regularity of the residual, $\Delta(A,f)$, by estimates on the terms in the integrand. Consider an arbitrary term, $(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})$, in the series, $\sum_{m=1}^{d}(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})$, in the integrand. Then
@@
from Holders inequality. Handling the estimates of $\lVert A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}$ and $\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$ separately, we have:
@@
$=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$ (44)
where $\alpha^{\mathbf{j}}_{\mathbf{k}}$ is the wavelet coefficient at the specified scale and location parameters, the second inequality comes from triangle inequality, the third inequality from the $L^{1}$ estimate on $A^{m}$- which is finite since $A^{m}(Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\in L^{\infty}$ by the maximum principle - and that wavelet coefficients of Holder functions can be characterized by exponential decay in the scale parameters [7, 8]. For $\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$,
@@
$=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$ (45)
where again, the last inequality comes from exponential decay in the wavelet scaling parameter and Holder functions. Combining the estimates on $\lVert A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}$ and $\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$ yields:
$\lVert A^{m}(\delta^{t}(f))\mathbf{v}^{m}\rVert_{L^{1}([0,1]^{d})}\leq C2^{-(j_{1}+j_{2}+\ldots,j_{d})2\alpha}$ (46)
Now we bound the approximation terms, $\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}$, for the mth term in the series under the integral for $\Delta(A,f)$ similarly. By Holder’s inequality,
$\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}\leq\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\rVert_{L^{1}([0,1]^{d})}\lVert\tilde{\mathbf{v}}^{m}\rVert_{L^{\infty}([0,1]^{d})}$ (47)
@@
$\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$ (48)
@@
$=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$ (49)
@@
$\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$ (50)
@@
$\left|\lVert(A^{m}(\delta^{t}(f))\mathbf{v}^{m}\rVert_{L^{1}([0,1]^{d})}-\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}\right|\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$ (51)
@@
$(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\tilde{\mathbf{v}}^{1}$ (52)
@@
$\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$ (53)
by application of Holder’s inequality, the property of $A\in C^{d}$, and the estimates on the wavelet coefficients of Holder functions previously discussed. So by the previous result, the first term in the series under the integrand of $\Delta(A,f)$ has the estimate:
$\left|\lVert(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\tilde{\mathbf{v}})\rVert_{L^{1}([0,1]^{d})}\right|\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$ (54)
@@
$=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$ (55)
@@
$\lVert A^{d}(\delta^{t}(f))\mathbf{v}^{d}\rVert_{L^{1}([0,1]^{d})}\leq C2^{-(j_{1}+j_{2}+\ldots,j_{d})2\alpha}$ (56)
@@
$=C2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$ (57)
Combining (56) and (57), the last term of the series in the integrand of $\Delta(A,f)$ has the estimate:
$\left|\lVert(A^{d}(\delta^{t}(f))\mathbf{v}^{d}\rVert_{L^{1}([0,1]^{d})}-\lVert A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d})\rVert_{L^{1}([0,1]^{d})}\right|\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$ (58)
So by (51), (54), and (58) we can express the residual, $\Delta(A,f)$, defined in (42), as a sum of piecewise constant functions:
$\Delta(A,f)=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}(\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\beta^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}\chi_{I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}}}(x_{1},x_{2},\ldots,x_{d})),$
$|\beta^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}|\leq C2^{-(j_{1}+j_{2}+\ldots j_{d})2\alpha}$ (59)
Now we characterize the regularity of $\Delta(A,f)$ with the following lemma:
Lemma 3.
The residual, $\Delta(A,f)=A(f)-\tilde{A}(f)$, defined as:
$\Delta(A,f)=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}(\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\beta^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}\chi_{I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}}}(x_{1},x_{2},\ldots,x_{d})),$
$|\beta^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}|\leq C2^{-(j_{1}+j_{2}+\ldots j_{d})2\alpha}([0,1]^{d})$ (60)
which is the sum of piecewise constant functions with bounded coefficients, is an element of $\Lambda^{2\alpha}$
@@
To show $\Delta(A,f)\in\Lambda^{2\alpha}([0,1]^{d})$, we need to show
$|\Delta(A,f)(\mathbf{x})-\Delta(A,f)(\mathbf{\tilde{x}})|\leq Cd_{1}(x_{1},\tilde{x}_{1})^{2\alpha}d_{2}(x_{2},\tilde{x}_{2})^{2\alpha}\ldots d_{d}(x_{d},\tilde{x}_{d})^{2\alpha}$ (61)
@@
Then summing over the scaling and location parameters in $\mathcal{U}$ for $\Delta(A,f)$, one has:

### v2

@@ 3. Preliminaries
$\mathbf{v}^{d-1}(f)\coloneq\frac{\partial^{2}}{\partial t_{d}\partial t_{d-1}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-2}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))+\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{d-1}\partial t_{d-2}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))$
(28) $+\ldots+\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{3}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{2}\partial t_{1}}(\delta^{t}(f))$
(29) $\mathbf{v}^{d}(f)\coloneq\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))$
@@
$A(f)=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}\int_{0}^{1}\int_{0}^{1}\ldots\int_{0}^{1}A^{\prime}(\delta^{t}(f))\mathbf{v}^{1}+A^{\prime\prime}(\delta^{t}(f))\mathbf{v}^{2}+\ldots+$
(30) $A^{m}(\delta^{t}(f))\mathbf{v}^{m}+\ldots+A^{d-1}(\delta^{t}(f))\mathbf{v}^{d-1}+A^{d}(\delta^{t}(f))\mathbf{v}^{d}dt_{1}dt_{2}\ldots dt_{d}$
We can now define the approximation, $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$, by replacing the sequence $(\mathbf{v}^{1},\mathbf{v}^{2}\ldots,\mathbf{v}^{d})$ with $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ - each element in the latter sequence is comprised with a wavelet coefficient in the direction of the partial derivative for an element in the former sequence.
(31) $\mathbf{\tilde{v}}^{1}(f)\coloneq Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}}(f)$
$\tilde{\mathbf{v}}^{2}(f)\coloneq Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{2}}(f)Q^{j_{1}}(f)+Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{3}}(f)Q^{j_{2}}Q^{j_{1}}(f)+\ldots$
(32) $+Q^{j_{d}}(f)Q^{j_{d-1}}Q^{j_{d-2}}\ldots Q^{j_{1}}(f)$
@@
$\tilde{\mathbf{v}}^{m}(f)\coloneq Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{m}}(f)Q^{j_{m-1}}(f)\ldots Q^{j_{1}}(f)$
@@
$\tilde{\mathbf{v}}^{d-1}(f)\coloneq Q^{j_{d}}Q^{j_{d-1}}(f)Q^{j_{d-2}}(f)\ldots Q^{j_{1}}(f)+Q^{j_{d}}(f)Q^{j_{d-1}}Q^{j_{d-2}}(f)\ldots Q^{j_{1}}(f)$
(34) $+\ldots+Q^{j_{d}}(f)Q^{j_{d-1}}(f))\ldots Q^{j_{3}}(f)Q^{j_{2}}Q^{j_{1}}(f)$
(35) $\tilde{\mathbf{v}}^{d}(f)\coloneq Q^{j_{d}}(f)Q^{j_{d-1}}(f)\ldots Q^{j_{1}}(f)$
Again, note the first, $\tilde{\mathbf{v}}^{1}$, and last, $\tilde{\mathbf{v}}^{d}$, operators have one term while the operators, $(\mathbf{v}^{2},\ldots,\mathbf{v}^{d-1})$, are comprised of a series each with d1d-1 terms. In this case, each term in the series contains the number of wavelet coefficients corresponding to the power of the operator (i.e. the term $\tilde{\mathbf{v}}^{m}$ has $m$ wavelet coefficients for each term in its series). Now the following linear combination of $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ is independent of $t$ and is used to construct the approximation, $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$:
$\tilde{A}_{(N_{i})_{i=0}^{d}}(f)\coloneq\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}+A^{\prime\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{2}+\ldots+A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m}$
@@
(36) $=\sum_{j_{d}=0_{d}}^{N_{d}}\sum_{j_{d-1}=0_{d-1}}^{N_{d-1}}\ldots\sum_{j_{1}=0_{1}}^{N_{1}}(\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta})$
Computing the residual, $\Delta_{(N_{i})_{i=1}^{d}}(A,f)\coloneq A(f)-\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$, yields:
@@
(37) $=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}\int_{0}^{1}\int_{0}^{1}\ldots\int_{0}^{1}\sum_{m=1}^{d}(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})dt_{1}dt_{2}\ldots dt_{d}$
We characterize the regularity of the residual, $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$, by estimates on the terms in the integrand. Consider an arbitrary term, $(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})$, in the series, $\sum_{m=1}^{d}(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})$, in the integrand. Then
@@
from Hölders inequality. Handling the estimates of $\lVert A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}$ and $\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$ separately, we have:
@@
(39) $=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$
where $\alpha^{\mathbf{j}}_{\mathbf{k}}$ is the wavelet coefficient at the specified scale and location parameters, the second inequality comes from triangle inequality, the third inequality from the $L^{1}$ estimate on $A^{m}$- which is finite since $A^{m}(Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\in L^{\infty}$ by the maximum principle - and that wavelet coefficients of Hölder functions can be characterized by exponential decay in the scale parameters [4, 3]. For $\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$,
@@
(40) $=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$
where again, the last inequality comes from exponential decay in the wavelet scaling parameter and Hölder functions. Combining the estimates on $\lVert A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}$ and $\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$ yields:
(41) $\lVert A^{m}(\delta^{t}(f))\mathbf{v}^{m}\rVert_{L^{1}([0,1]^{d})}\leq C2^{-(j_{1}+j_{2}+\ldots,j_{d})2\alpha}$
Now we bound the approximation terms, $\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}$, for the mth term in the series under the integral for $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ similarly. By Hölder’s inequality,
(42) $\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}\leq\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\rVert_{L^{1}([0,1]^{d})}\lVert\tilde{\mathbf{v}}^{m}\rVert_{L^{\infty}([0,1]^{d})}$
@@
(43) $\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$
@@
(44) $=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$
@@
(45) $\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$
@@
(46) $\left|\lVert(A^{m}(\delta^{t}(f))\mathbf{v}^{m}\rVert_{L^{1}([0,1]^{d})}-\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}\right|\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$
@@
(47) $(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\tilde{\mathbf{v}}^{1}$
@@
(48) $\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$
by application of Hölder’s inequality, the property of $A\in C^{d}$, and the estimates on the wavelet coefficients of Hölder functions previously discussed. So by the previous result, the first term in the series under the integrand of $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ has the estimate:
(49) $\left|\lVert(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\tilde{\mathbf{v}})\rVert_{L^{1}([0,1]^{d})}\right|\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$
@@
(50) $=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$
@@
(51) $\lVert A^{d}(\delta^{t}(f))\mathbf{v}^{d}\rVert_{L^{1}([0,1]^{d})}\leq C2^{-(j_{1}+j_{2}+\ldots,j_{d})2\alpha}$
@@
(52) $=C2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$
Combining (51) and (52), the last term of the series in the integrand of $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ has the estimate:
(53) $\left|\lVert(A^{d}(\delta^{t}(f))\mathbf{v}^{d}\rVert_{L^{1}([0,1]^{d})}-\lVert A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d})\rVert_{L^{1}([0,1]^{d})}\right|\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$
So by (46), (49), and (53) we can express the residual, $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$, defined in (37), as a sum of piecewise constant functions:
$\Delta_{(N_{i})_{i=1}^{d}}(A,f)=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}(\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\eta^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}\chi_{I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}}}(x_{1},x_{2},\ldots,x_{d})),$
(54) $|\beta^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}|\leq C2^{-(j_{1}+j_{2}+\ldots j_{d})2\alpha}$
Now we characterize the regularity of $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ with the following lemma:
Lemma 4.5.
The residual, $\Delta_{(N_{i})_{i=1}^{d}}(A,f)=A(f)-\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$, can be writen as:
$\Delta_{(N_{i})_{i=1}^{d}}(A,f)=\sum_{j_{1},\ldots,j_{d}=0}^{N_{1},\ldots,N_{d}}(\sum_{k_{1},\ldots,k_{d}=1}^{2^{j_{1}},\ldots,2^{j_{d}}}\eta^{(j_{i})_{i=1}^{d}}_{(k_{i})_{i=1}^{d}}\chi_{I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}}}(x_{1},x_{2},\ldots,x_{d})),$
(55) $|\eta^{(j_{i})_{i=1}^{d}}_{(k_{i})_{i=1}^{d}}|\leq C2^{-(j_{1}+j_{2}+\ldots j_{d})2\alpha}$
and is an element of $\Lambda_{2\alpha}([0,1]^{d})$
@@
To show $\Delta_{(N_{i})_{i=1}^{d}}(A,f)\in\Lambda_{2\alpha}([0,1]^{d})$, we need to show
(56) $|\Delta_{(N_{i})_{i=1}^{d}}(A,f)(\mathbf{x})-\Delta_{(N_{i})_{i=1}^{d}}(A,f)(\mathbf{\tilde{x}})|\leq Cd_{d}(x^{\prime}_{1},x_{1})^{2\alpha}d_{d}(x^{\prime}_{2},x_{2})^{2\alpha}\ldots d_{d}(x^{\prime}_{d},x_{d})^{2\alpha}$
@@
Then summing over the scaling and location parameters in $\mathcal{U}$ for $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$, one has: